Elen wishes to mix candy worth 2.15/ pound with candy worth 3.92/ pound to produce 30 pounds of mixture worth 2.80/pound. How many pounds of the more expensive candy should he use?
Found 2 solutions by checkley77, stanbon:
Answer by checkley77(12844)

3.92x+2.15(30-x)=2.80*30
3.92x+64.5-2.15x=84
1.77x=84-64.5
1.77x=19.5
x=19.5/1.77
x=11.017 lbs. of $3.92 candy is used.
30-11.017=18.983 lbs. of the cheap candy.
Proof:
3.92*11.017+2.15*18.983=84
43.1866+40.8136=84
84~84

Answer by stanbon(75887)

Elen wishes to mix candy worth 2.15/ pound with candy worth 3.92/ pound to produce 30 pounds of mixture worth 2.80/pound. How many pounds of the more expensive candy should he use?
-------------------------------
Value Equation:
value + value = value
2.15x + 3.92(30-x) = 2.80*30
Multiply thru by 100 to get:
215x + 392*30 - 392x = 280*30
-177x = -112*30
x = 18.98 lbs.(amt. of 2.15 candy)
30 - 18.98 = 11.02 lbs (amt. of 3.92 candy)
===============================================
Cheers,
Stan H.